# FlowAlign: Trajectory-Regularized, Inversion-Free Flow-based Image Editing

### v1

@@ Appendix A Optimal Control Formulation to Prove Proposition 1
$\dot{\bm{x}}_{t}=-\frac{2\lambda}{(1+\lambda)^{2}}{\bm{r}}+\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})$ (30)
where the constant resolution ${\bm{r}}$ is given by
@@
The optimal ${\bm{u}}_{t}^{*}$ that minimizes the Hamitonian is then given by
@@
$H({\bm{x}}_{t},{\bm{u}}^{*}_{t},{\bm{p}}_{t},t):=$ $\frac{1}{2(1+\lambda)^{2}}\left(\|{\bm{p}}_{t}-({\bm{a}}_{t}-{\bm{b}}_{t})\|^{2}+\|{\bm{p}}_{t}-({\bm{b}}_{t}-{\bm{a}}_{t})\|^{2}\right)$
@@
Then, the minimum principle Basar et al. [2020], Fleming and Rishel [2012] informs that the optimal pair $({\bm{x}}_{t},{\bm{p}}_{t})$ should satisfy the following:
$\dot{\bm{x}}_{t}$ $=\frac{\partial H({\bm{x}}_{t},{\bm{u}}^{*}_{t},{\bm{p}}_{t},t)}{\partial{\bm{p}}}=-\frac{2\lambda}{(1+\lambda)^{2}}{\bm{p}}_{t}+\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})$ (35)
@@
${\bm{x}}_{0}={\bm{x}}_{start},$ $\quad{\bm{p}}_{1}=\left.\frac{\partial m({\bm{x}}_{t},t)}{\partial{\bm{x}}_{t}}\right|_{t=1}=\eta({\bm{x}}_{1}-{\bm{x}}_{end})$ (37)
From Eq. (36), we can see that ${\bm{p}}_{t}$ is time-invariant constant, i.e. ${\bm{p}}_{t}={\bm{r}}$ for all $t\in[0,1]$. Combining this with Eq. (35), we have
${\bm{x}}_{1}$ $={\bm{x}}_{0}+\int_{0}^{1}-\frac{2\lambda}{(1+\lambda)^{2}}{\bm{r}}+\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})dt$ (38)
$=-\frac{2\lambda}{(1+\lambda)^{2}}{\bm{r}}+\hat{\bm{x}}_{1}$ (39)
@@
By plugging this in Eq. (37) with ${\bm{p}}_{1}={\bm{r}}$, we have
${\bm{r}}=\eta\left(\hat{\bm{x}}_{1}-{\bm{x}}_{end}-\frac{2\lambda}{(1+\lambda)^{2}}{\bm{r}}\right)$$\Rightarrow$ ${\bm{r}}=\frac{\eta}{1+\frac{2\lambda\eta}{(1+\lambda)^{2}}}\left(\hat{\bm{x}}_{1}-{\bm{x}}_{end}\right)$ (41)
@@
This leads to the modified form of the residual term:
@@ Appendix B Implementation details
While various text-to-image generative models exist, we leverage a flow-based model defined in latent space, specifically using Stable Diffusion 3.0 (medium) Esser et al. [2024] provided by the diffusers package. Given the significantly improved generative and text alignment performance of this backbone compared to earlier versions, we use the same backbone model for all baselines to establish a fair comparison. For the time discretizations, we set the shift coefficient to 3.0, which is a default option of the Stable Diffusion 3.0
@@
    DDIB Su et al. [2023] : DDIB involves an inversion process followed by a sampling process. For inversion, we adopt the backward flow ODE. Regarding classifier-free guidance (CFG), we use only the null-text embedding during inversion and both target text and null-text embeddings during the sampling. This choice is motivated by the instability observed when applying standard CFG (i.e., using the source text and null-text embeddings) during inversion. To ensure a fair comparison under similar computational cost, we set the number of ODE timesteps to 17 for both the inversion and sampling processes.
@@
    SDEdit Meng et al. [2021] : SDEdit requires specifying the initial SNR (i.e. timestep), and its performance can vary significantly depending on this choice. In this work, our main focus is to address the limitations of inversion-free editing methods. To fairly demonstrate the effectiveness of our approach in comparison to alternative methods that also can mitigate this issue, we adopt the same initial SNR setting as FlowEdit Kulikov et al. [2024], determined by the starting timestep of the ODE. Specifically, we use the 18th timestep as the starting point in our experiments.
@@
    RF-inversion Rout et al. [2025a] : RF-Inversion introduces an optimal-control-based guidance mechanism that ensures the inverted representation aligns with a target terminal state, resuling in a sampling process that is more likely under a predefined terminal distribution. We follow the official implementation, setting $\gamma=0.5$, $\eta=0.9$, the starting time $s=0$, and the stopping time $\tau=0.25$. RF-Inversion uses only the null-text embedding, while both the target text and null-text embeddings are used during the sampling phase.
@@
    FlowEdit Kulikov et al. [2024] : We follow the official implementation of FlowEdit, setting the CFG scale to 3.0 for the source direction and 13.5 for the target direction. Additionally, we solve the flow ODE starting from the 18th timestep out of 50, resulting in 33 ODE timesteps.
@@
For the quantitative comparison, we evaluate following metrics using the official evaluation code  [^3] https://github.com/cure-lab/PnPInversion/tree/main/evaluation from PIEBench Ju et al. [2024]:
@@
    Background PSNR : PIEbench Ju et al. [2024] provides masks that cover the object to be edited. Accordingly, we compute the PSNR by excluding the masked region, resulting in the background PSNR.
@@
    Background LPIPS : We measure the LPIPS Blau and Michaeli [2018], which is defined as distance between feature maps of pre-trained VGG network, by excluding the masked region.
@@
    Background SSIM : We compute the structural similarity Wang et al. [2004] by excluding the masked region.
@@
    CLIP-score : We report the similarity between features embedded by pre-trained CLIP Radford et al. [2021] [^4] We use CLIP ViT-base-patch16. image encoder and text encoder. For the CLIP score within the edited region, we apply it only to the masked area.
@@ Appendix C Human preference test protocol
To evaluate the quality of image editing, we conduct a human preference study in the form of an AB-test. Specifically, we randomly sample 100 images from the 700 validation samples of PIEBench Ju et al. [2024]. The study follows the protocol below for each participant:
@@ E.1 Quantitative and qualitative results
DDIB Su et al. [2023] 0.106 14.43 0.262 0.037 0.746 30.68 26.88
SDEdit Meng et al. [2021] 0.036 22.57 0.119 0.008 0.747 24.56 21.95
RF-Inv Rout et al. [2025a] 0.103 15.82 0.291 0.028 0.663 33.35 29.60
FlowEdit Kulikov et al. [2024] 0.036 23.02 0.082 0.007 0.842 25.98 22.81
@@ F.1 Further Applications - Video Editing
FlowAlign is fundamentally a text-based image editing method built on an image flow-based model. In contrast, video editing typically relies on generative models trained on video datasets with temporal attention mechanisms to enhance temporal consistency Jeong et al. [2024b], Liu et al. [2024], Jeong et al. [2024a], Park et al. [2025]. However, video editing can also be viewed as a sequence of image editing tasks. Extending the applicable range of image editing methods to video could offer practical benefits, such as reduced training and inference costs.
While the proposed method does not explicitly enforce temporal consistency, we can still apply it independently to each video frame. For this experiment, we use the DAVIS dataset Caelles et al. [2019] and extract the source text prompt using LLaVA Liu et al. [2023a], conditioned on the middle frame of each video. For the hyperparameter $\lambda$, we use a constant of 0.01, consistent with the main experiments. For $\omega$, we vary it over [5.0, 7.5, 10.0, 13.5] and qualitatively select the best-performing value.
@@ F.2 Further Applications - Gaussian Splatting Editing
We perform Gaussian Splatting optimization for $3,000$ iterations using a classifier-free guidance (CFG) weight $\omega$ set to either 70 or 100. We empirically set $\lambda_{t}=30\sigma_{t}^{2}$ where $\sigma_{t}=t$, which monotonically decreases toward zero as $t$ decreases from 0.9 to 0.1 as in the image editing case. Thus, flow-matching guidance is emphasized during the early Gaussian updates. We set the step size $\eta_{t}=1+30\sigma_{t}^{2}$. This setting consistently achieves a favorable balance between semantic fidelity and structural integrity. Experiments are conducted on real-world scenes using datasets from IN2N Haque et al. [2023] and PDS Koo et al. [2024]. For comparison, we benchmark against existing distillation methods, including SDS Poole et al. [2022], DDS Hertz et al. [2023], and PDS Koo et al. [2024], as well as the non-distillation baseline IGS2GS Vachha and Haque [2024]. For baselines using distillation methods we used Stable Diffusion 3.0 and for IGS2GS Vachha and Haque [2024], we use UltraEdit Zhao et al. [2024] in replacement of Instruct-Nerf2Nerf  Haque et al. [2023] for flow-based models. One can find the pseudocode for the 3D editing with FlowAlign in Algorithm 2, and additional results on figure15 and figure 15.
@@ F.3 Further Variants - FLUX
Our main experiments are conducted using by leveraging pre-trained flow model, Stable Diffusion 3.0. Since the proposed method regulates the sampling ODE via a flow matching cost, it is also compatible with other flow-based models, such as FLUX. Following the setup in FlowEdit Kulikov et al. [2024], we additionally evaluate our method using FLUX as the backbone model. We incorporate the source consistency term derived from optimal control with flow matching regularization, as described in Algorithm 3. As FLUX.1-dev is a guidance-distilled model that directly takes the CFG scale as input, there is no need to apply the modified CFG strategy used in Stable Diffusion 3.0 case. Therefore, we simply provide the CFG scale to both ${\bm{v}}({\bm{p}}_{t})$ and ${\bm{v}}({\bm{q}}_{t})$, following the approach used in FlowEdit. Figure 16 illustrates edited results generated by the proposed method implemented with FLUX.1-dev. Similar to the results with Stable Diffusion 3.0, the outputs effectively reflect the intended editing direction specified by the text prompts while preserving source structures. These results imply that the proposed method is broadly applicable to flow-based models for image editing.

### v2

@@ Appendix A Optimal Control Formulation to Prove Proposition 1
$\dot{\bm{x}}_{t}=-\frac{1}{1+\lambda}{\bm{p}}_{t}+\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})$ (30)
where ${\bm{p}}_{t}$ is given by
@@
The optimal control ${\bm{u}}_{t}^{*}$ that minimizes the Hamitonian is then given by
@@
$H({\bm{x}}_{t},{\bm{u}}^{*}_{t},{\bm{p}}_{t},t):=$ $\frac{1}{2(1+\lambda)^{2}}\left(\|{\bm{p}}_{t}-\lambda({\bm{a}}_{t}-{\bm{b}}_{t})\|^{2}+\lambda\|{\bm{p}}_{t}-({\bm{b}}_{t}-{\bm{a}}_{t})\|^{2}\right)$
@@
Then, the minimum principle [1, 6] informs that the optimal pair $({\bm{x}}_{t},{\bm{p}}_{t})$ should satisfy the following:
$\dot{\bm{x}}_{t}$ $=\frac{\partial H({\bm{x}}_{t},{\bm{u}}^{*}_{t},{\bm{p}}_{t},t)}{\partial{\bm{p}}}$
$=\frac{1}{(1+\lambda)^{2}}({\bm{p}}_{t}-\lambda({\bm{a}}_{t}-{\bm{b}}_{t}))+\frac{\lambda}{(1+\lambda)^{2}}({\bm{p}}_{t}-({\bm{b}}_{t}-{\bm{a}}_{t}))+\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})-\frac{2}{1+\lambda}{\bm{p}}_{t}$
$=-\frac{1}{1+\lambda}{\bm{p}}_{t}+\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})$ (35)
@@
${\bm{x}}_{t_{0}}={\bm{x}}_{start},$ $\quad{\bm{p}}_{1}=\left.\frac{\partial m({\bm{x}}_{t},t)}{\partial{\bm{x}}_{t}}\right|_{t=1}=\eta({\bm{x}}_{1}-{\bm{x}}_{end})$ (37)
From Eq. (36), we can see that ${\bm{p}}_{t}$ is time-invariant constant, i.e. ${\bm{p}}_{t}={\bm{p}}$ for all $t\in[0,1]$. Combining this with Eq. (35), we have
${\bm{x}}_{1}$ $={\bm{x}}_{t_{0}}+\int_{t_{0}}^{1}-\frac{1}{(1+\lambda)}{\bm{p}}+\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})dt$ (38)
$=-\frac{1-t_{0}}{1+\lambda}{\bm{p}}+{\bm{x}}_{t_{0}}+{\bm{v}}_{t_{0}}$ (39)
@@
By plugging this in Eq. (37) with ${\bm{p}}_{1}={\bm{p}}$, we have
${\bm{p}}=\eta\left({\bm{x}}_{t_{0}}+{\bm{v}}_{t_{0}}-{\bm{x}}_{end}-\frac{(1-t_{0})}{(1+\lambda)}{\bm{p}}\right)$$\Rightarrow$ ${\bm{p}}=\frac{\eta}{1+\frac{\eta(1-t_{0})}{(1+\lambda)}}\left({\bm{x}}_{t_{0}}+{\bm{v}}_{t_{0}}-{\bm{x}}_{end}\right)$
Therefore, the optimal control ${\bm{u}}_{t}^{*}$ is given by
${\bm{u}}_{t}^{*}$ $=\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})-\frac{1}{1+\lambda}{\bm{p}}_{t}$ (41)
$=\frac{1}{1+\lambda}({\bm{a}}_{t}+\lambda{\bm{b}}_{t})-\frac{\eta}{(1+\lambda)+\eta(1-t)}\left({\bm{x}}_{t}+{\bm{v}}_{t}-{\bm{x}}_{end}\right)$ (42)
@@
This leads to the following:
@@ Appendix B Implementation details
While various text-to-image generative models exist, we leverage a flow-based model defined in latent space, specifically using Stable Diffusion 3.0 (medium) [5] provided by the diffusers package. Given the significantly improved generative and text alignment performance of this backbone compared to earlier versions, we use the same backbone model for all baselines to establish a fair comparison. For the time discretizations, we set the shift coefficient to 3.0, which is a default option of the Stable Diffusion 3.0
@@
    DDIB [32] : DDIB involves an inversion process followed by a sampling process. For inversion, we adopt the backward flow ODE. Regarding classifier-free guidance (CFG), we use only the null-text embedding during inversion and both target text and null-text embeddings during the sampling. This choice is motivated by the instability observed when applying standard CFG (i.e., using the source text and null-text embeddings) during inversion. To ensure a fair comparison under similar computational cost, we set the number of ODE timesteps to 17 for both the inversion and sampling processes.
@@
    SDEdit [21] : SDEdit requires specifying the initial SNR (i.e. timestep), and its performance can vary significantly depending on this choice. In this work, our main focus is to address the limitations of inversion-free editing methods. To fairly demonstrate the effectiveness of our approach in comparison to alternative methods that also can mitigate this issue, we adopt the same initial SNR setting as FlowEdit [16], determined by the starting timestep of the ODE. Specifically, we use the 18th timestep as the starting point in our experiments.
@@
    RF-inversion [28] : RF-Inversion introduces an optimal-control-based guidance mechanism that ensures the inverted representation aligns with a target terminal state, resuling in a sampling process that is more likely under a predefined terminal distribution. We follow the official implementation, setting $\gamma=0.5$, $\eta=0.9$, the starting time $s=0$, and the stopping time $\tau=0.25$. RF-Inversion uses only the null-text embedding, while both the target text and null-text embeddings are used during the sampling phase.
@@
    FlowEdit [16] : We follow the official implementation of FlowEdit, setting the CFG scale to 3.0 for the source direction and 13.5 for the target direction. Additionally, we solve the flow ODE starting from the 18th timestep out of 50, resulting in 33 ODE timesteps.
@@
For the quantitative comparison, we evaluate following metrics using the official evaluation code  [^3] https://github.com/cure-lab/PnPInversion/tree/main/evaluation from PIEBench [13]:
@@
    Background PSNR : PIEbench [13] provides masks that cover the object to be edited. Accordingly, we compute the PSNR by excluding the masked region, resulting in the background PSNR.
@@
    Background LPIPS : We measure the LPIPS [2], which is defined as distance between feature maps of pre-trained VGG network, by excluding the masked region.
@@
    Background SSIM : We compute the structural similarity [35] by excluding the masked region.
@@
    CLIP-score : We report the similarity between features embedded by pre-trained CLIP [27] [^4] We use CLIP ViT-base-patch16. image encoder and text encoder. For the CLIP score within the edited region, we apply it only to the masked area.
@@ Appendix C Human preference test protocol
To evaluate the quality of image editing, we conduct a human preference study in the form of an AB-test. Specifically, we randomly sample 100 images from the 700 validation samples of PIEBench [13]. The study follows the protocol below for each participant:
@@ E.1 Quantitative and qualitative results
DDIB [32] 0.106 14.43 0.262 0.037 0.746 30.68 26.88
SDEdit [21] 0.036 22.57 0.119 0.008 0.747 24.56 21.95
RF-Inv [28] 0.103 15.82 0.291 0.028 0.663 33.35 29.60
FlowEdit [16] 0.036 23.02 0.082 0.007 0.842 25.98 22.81
@@ F.1 Further Applications - Video Editing
FlowAlign is fundamentally a text-based image editing method built on an image flow-based model. In contrast, video editing typically relies on generative models trained on video datasets with temporal attention mechanisms to enhance temporal consistency [11, 19, 10, 24]. However, video editing can also be viewed as a sequence of image editing tasks. Extending the applicable range of image editing methods to video could offer practical benefits, such as reduced training and inference costs.
While the proposed method does not explicitly enforce temporal consistency, we can still apply it independently to each video frame. For this experiment, we use the DAVIS dataset [4] and extract the source text prompt using LLaVA [18], conditioned on the middle frame of each video. For the hyperparameter $\lambda$, we use a constant of 0.01, consistent with the main experiments. For $\omega$, we vary it over [5.0, 7.5, 10.0, 13.5] and qualitatively select the best-performing value.
@@ F.2 Further Applications - Gaussian Splatting Editing
We perform Gaussian Splatting optimization for $3,000$ iterations using a classifier-free guidance (CFG) weight $\omega$ set to either 70 or 100. We empirically set $\lambda_{t}=30\sigma_{t}^{2}$ where $\sigma_{t}=t$, which monotonically decreases toward zero as $t$ decreases from 0.9 to 0.1 as in the image editing case. Thus, flow-matching guidance is emphasized during the early Gaussian updates. We set the step size $\eta_{t}=1+30\sigma_{t}^{2}$. This setting consistently achieves a favorable balance between semantic fidelity and structural integrity. Experiments are conducted on real-world scenes using datasets from IN2N [7] and PDS [15]. For comparison, we benchmark against existing distillation methods, including SDS [26], DDS [8], and PDS [15], as well as the non-distillation baseline IGS2GS [34]. For baselines using distillation methods we used Stable Diffusion 3.0 and for IGS2GS [34], we use UltraEdit [36] in replacement of Instruct-Nerf2Nerf  [7] for flow-based models. One can find the pseudocode for the 3D editing with FlowAlign in Algorithm 2, and additional results on figure15 and figure 15.
@@ F.3 Further Variants - FLUX
Our main experiments are conducted using by leveraging pre-trained flow model, Stable Diffusion 3.0. Since the proposed method regulates the sampling ODE via a flow matching cost, it is also compatible with other flow-based models, such as FLUX. Following the setup in FlowEdit [16], we additionally evaluate our method using FLUX as the backbone model. We incorporate the source consistency term derived from optimal control with flow matching regularization, as described in Algorithm 3. As FLUX.1-dev is a guidance-distilled model that directly takes the CFG scale as input, there is no need to apply the modified CFG strategy used in Stable Diffusion 3.0 case. Therefore, we simply provide the CFG scale to both ${\bm{v}}({\bm{p}}_{t})$ and ${\bm{v}}({\bm{q}}_{t})$, following the approach used in FlowEdit. Figure 16 illustrates edited results generated by the proposed method implemented with FLUX.1-dev. Similar to the results with Stable Diffusion 3.0, the outputs effectively reflect the intended editing direction specified by the text prompts while preserving source structures. These results imply that the proposed method is broadly applicable to flow-based models for image editing.